# Addressing Blind Guessing: Calibration of Selection Bias in Multiple-Choice Question Answering by Video Language Models

### v1

@@ 4.1 Evaluation Criteria
Our framework both detects bias and evaluates performance. Bias metrics assess fairness by evaluating the model’s decision making on consistency in treating all answer options, while performance metrics evaluate accuracy and reliability.
We monitor the standard deviations of the F1 score (F1_std) and the recall (Recall_std). High F1_std indicates inconsistent performance across options, signaling bias. High Recall_std shows the model is better at correctly identifying answers for certain options over others.
For option probabilities, we use the standard deviation of the symmetric Jensen-Shannon distance (JS_std) to detect inconsistencies in how the predicted probability distributions differ from the true distributions. This metric, bounded between 0 and 1, is more interpretable than Kullback-Leibler (KL) divergence.
@@ 4.2 Debiasing Approach
Every MCQA benchmark comprises essential components that make the answering task feasible. In the case of video-MCQA, these components include the video, the question, and the answer options. We construct decompositions where the task becomes impossible by design, as one of the key components is removed. This allows us to examine the projections of model’s behavioral bias on each decomposition plane as shown in Figure 5. In a fair and unbiased model, projecting the dataset by omitting one of the key components would result in a uniform distribution. However, as shown in Section 3, this is far from the case. In the decomposition planes, the bias is more clearly manifested, and since we expect uniformity, the bias becomes easier to detect and mitigate.
@@ 4.2.1 Decomposition
Consider a task $T$ that we wish to decompose into $n$ non-overlapping components $C_{1},C_{2},\dots,C_{n}$. This decomposition can be expressed as:
$T=\bigcup_{i=1}^{n}C_{i},\quad\text{where}\quad C_{i}\cap C_{j}=\emptyset\quad\text{for all}\quad i\neq j$
The task is considered ill-defined if it lacks one of the key components. Additionally, there is a well-defining property, or correctness, $P(T)$ of the task. If any component $C_{i}$ is removed, the property $P(T)$ becomes ill defined. We assume that all tasks are initially well-defined.
@@
For an ill-defined task, a model that is both reasonable and unbiased should uniformly select any option ID. Formally, if $P(T)$ is ‘ill-defined’: $M(T)=\text{RandomChoice}(\{ID_{opt}\})$.
The key insight is that if the model has bias on a given dataset, this bias is likely to be more ubiquitous, hence evident, in the ill-defined versions of the tasks. Thus, the model bias can be decomposed as follows: $Bias_{M}\supseteq\bigcup Bias_{M}|_{\hat{D}}$, where $\hat{D}$ represents an ill-defined copy of the dataset $D$.
Let $A$ be an attack on the well-defined task that makes this task defined insufficiently. Subsequently, there are three types of attacks that make the task ill-defined: $A|_{v=0}$ for all zeroed frames in the video, $A|_{q=0}$ for all questions as an empty string, and $A|_{o=0}$ for all options as IDs only.
We investigate the importance of the following components of the bias: $Bias_{M}|_{A|_{v=0}(D)}$, $Bias_{M}|_{A|_{q=0}(D)}$, $Bias_{M}|_{A|_{o=0}(D)}$, and how efficiently we can handle the general bias by balancing through these components.
@@ 4.2.2 Debiasing
First, we estimate the model bias on the dataset after attack A. We derive the $P_{p}$ knowing that $P_{d}$ is uniformly distributed from Equation  2:
@@
To estimate the bias under $A|_{v=0}$, $A|_{q=0}$ and $A|_{o=0}$ attacks, we first compute $P_{p}(d_{i}|A|_{v=0}(T))$, $P_{p}(d_{i}|A|_{q=0}(T))$ and $P_{p}(d_{i}|A|_{o=0})$, and then the prior estimation, respectively: $\tilde{P}_{p}(d_{i}|T)=softmax[P_{p}(d_{i}|A|_{v=0}(T))+P_{p}(d_{i}|A|_{q=0}(T))+P_{p}(d_{i}|A|_{o=0})]$.
@@
To gain the optimal weights, we use a 5-fold cross-validation procedure. In each fold, we optimize weights $\{w_{i}\}$ using the COBYLA algorithm Powell (1994), subject to the constraints $0\leq w_{i}\leq 1$ or $|w_{i}|\leq 1$. While we minimize the standard deviation of Recall_std across the options on the test set, we monitor the same bias metrics on the validation set. After completing the cross-validation we average the estimated global priors across the folds and apply this averaged global prior.
Both algorithms, BOLD and Weighted_BOLD, are outlined in Appendix D.
@@ 5 Results
Both BOLD and Weighted_BOLD consistently improve performance metrics for Video-LLaMA and Video-LLaVA. SeViLA generally exhibits greater resilience to bias calibration, showing minimal effect on performance on NExT-QA and marginal deterioration on STAR. This may be attributed to the model’s training for statistically close to uniform distribution of results over a fixed number of options. Despite decreases in Accuracy and F1 Mean on STAR, the reduction in Recall_std across all datasets for this model suggests that its debiased responses are more truthful, reflecting a fairer distribution of correct answers across options.
@@ 6 Conclusion
We performed the first comprehensive analysis on modified MCQA datasets to localize selection bias in specific options across VLMs of various architectures. Our findings reveal that while all models exhibit common behavioral patterns in their answer distributions, each model also has unique characteristics both across and within datasets. Despite these differences, the bias vector consistently manifests more prominently in the projections on the decomposed planes of video-question, video-answer, and question-answer interactions with one MCQA key component removed.
Using bias metrics to detect inconsistencies between the predicted and true probability distributions of the answers and calculating the prior bias, we proposed an effective method to mitigate this bias. Using a unique global prior for each dataset we reinforced the robustness and generalizability. Our BOLD bias removal approach not only enhances fairness, but also proves effective in improving performance metrics. Using weights in Weighted_BOLD further optimizes the performance.
This easily scalable, post-processing, automatic, and unsupervised method allows for a sharper focus on the true reasoning capabilities of the models. Our study contributes to making MCQA benchmarks more transparent and robust against the ordering of options, thereby delivering fairer and more meaningful evaluation results for VLMs.
Furthermore, elevating bias suppression to an abstract level, where tasks are systematically decomposed and key components are removed to identify where bias is most pronounced, can extend our approach to address other bias-related challenges in machine learning models.
@@ 7 Limitations
A primary limitation of our approach is the assumption that models rely on indirect cues when lacking sufficient direct information to determine an answer. Our conclusion—that bias can be decomposed across empty videos, questions, and answers—is based on empirical observations where bias is most pronounced in these settings. However, this decomposition may not be exhaustive; there could be more optimal directions in the latent space of decomposed priors. For instance, models often perform markedly different results compared to other settings when all answers are correct. A supervised debiasing method using true labels, such as applying COBYLA optimization directly, might also yield better results. Nonetheless, our goal was to find an easy-to-implement unsupervised solution without using correct answers.
We also assume the validity of the uniform distribution: Jensen-Shannon Divergence operates under the assumption that, in the absence of prior information, the distribution should be uniform. If the test MCQA data or task suggest an uneven distribution of answer options, this may lead to underestimating the significance of certain factors.
@@ 8 Ethics Statement
We based our experiments on existing datasets, ensuring compliance with their respective copyrights. Our research uses open-source VLMs, which, while offering accessibility and flexibility, also carry the inherent risks associated with open-ended text generation. Upon acceptance of the paper, we will release our code and dataset to the public for research and reproducibility.
@@ Acknowledgements
The first author thanks Amazon Alexa for supporting her research through a generous donation. We are deeply grateful to Prof. Ravi Shekhar and Giovanni Bonetta for their invaluable contributions at different stages of the project. In particular, we extend special thanks to Prof. Ravi Shekhar for developing the data preparation scripts that handled the movement of correct answers across options and the stratification of Perception Test, as well as for his consistent guidance and insightful consultations throughout the project.
@@ A.1 NExT-QA and NExT-GQA
The NExT-QA dataset, released in 2021, comprises videos sourced from the VidOR dataset Shang et al. (2019), with an average length of 35.73 seconds. The test split primarily contains causal questions (52.57%) addressing why and how, temporal reasoning questions (31.03%) about preceeding, following, or concurrent actions, and descriptive tasks (16.04%). Videos are manually annotated with questions and correct answers. Distractors are selected from 50 candidates based on a cosine similarity of less than 0.9 to the correct answer, ensuring they are close but not identical. This approach sometimes allows questions to be meaningfully answered without viewing the video (see Figures  68).
@@
NExT-GQA, released in 2024 as an extension of NExT-QA, focuses on evaluating the localization abilities of VLMs. It includes manually annotated timestamp labels that are key for answering the questions. Descriptive questions are excluded because they pertain to global video content or have answers available throughout the video with no specific temporal grounding. We also exclude QA pairs with multiple annotated time intervals. Consequently, the test set of NExT-GQA comprises 56.69% causal questions and 43.31% temporal questions.
@@ A.2 STAR
Situation hypergraphs are used to extract abstract representations from the videos. Multiple types of questions are then designed using functional programs to cover different levels of difficulty in situated reasoning. There are four types of questions focusing on occurred facts, temporal order, future probability, and feasibility in specific situations:
@@ A.3 Perception Test
The validation split we use contains approximately 15% closed questions, which poses an additional challenge for models that tend to exhibit affirmation bias Pezeshkpour and Hruschka (2023). Notably, for closed questions in this dataset, Video-LLaMA and Video-LLaVA often respond with "yes" or "no" rather than selecting an answer option.
@@ A.4 Video-MME
Video-MME was developed to address the lack of diversity in video MCQA benchmarks. It consists of 900 YouTube videos ranging from 11 seconds to 1 hour in length. The dataset includes 12 task types covering questions on temporal and spatial perception, object and action reasoning, OCR, and counting problems. All QA pairs and distractors are manually created with three questions per video. Additionally, the dataset features videos in languages other than English.
@@ B.1 Video-LLaMA
To address this issue, we experimented with various prompts, temperature settings, and token limits, and developed the following strategy:
@@ B.2 Video-LLaVA
Video-LLaVA effectively delivers concise answers as just a specific options based on a short, general prompt. However, in some cases, the model outputs as an answer option only "A" or responds with "yes" or "no" to a closed question.

### v2

@@ 4.1 Evaluation Criteria
Our framework both detects bias and evaluates performance. Bias metrics assess model’s consistency in treating all answer options fairly, while performance metrics evaluate accuracy and reliability.
We monitor the standard deviations of the F1 score (F1_std) and the recall (Recall_std). A high F1_std indicates inconsistent performance across options, signaling bias. A high Recall_std suggests the model is better at identifying correct answers for certain options over others.
For option probabilities, we use the standard deviation of the symmetric Jensen-Shannon distance (JS_std) to detect inconsistencies between the predicted and the true probability distributions. This metric, bounded between 0 and 1, is more interpretable than Kullback-Leibler (KL) divergence.
@@ 4.2 Debiasing Approach
Every MCQA benchmark comprises essential components that make the answering task feasible. In the case of video-MCQA, these components include the video, the question, and the answer options. We construct decompositions where the task becomes impossible by design, as one of the key components is removed. This allows us to examine the projections of model’s behavioral bias on each decomposition plane as intuitively shown in Figure 5. In a fair and unbiased model, projecting the dataset by omitting one of the key components would result in a uniform distribution. However, as shown in Section 3, this is far from the case. In the decomposition planes, the bias is more clearly manifested, and since we expect uniformity, the bias becomes easier to detect and mitigate.
@@ 4.2.1 Decomposition
Let a Multiple-Choice Question Answering task $T$ be well-defined if there is one and only one correct option to answer the question based on the provided context. Otherwise, the $T$ is ill-defined. We wish to decompose $T$ into the following independent components: video context, question, answer options. Omitting any of these components makes $T$ ill-defined. In general, there could be $n$ independent components $C_{1},C_{2},\dots,C_{n}$. Let $P(T)$ be a well-defining property (correctness) of $T$ such that $P(T)=1$ if $T$ is well-defined, otherwise - $P(T)=0$. We assume that all tasks are initially well-defined.
@@
For an ill-defined task, a model that is both reasonable and unbiased should uniformly select any option ID. Formally, if $P(T)==0$: $M(T)=\text{RandomChoice}(\{ID_{opt}\})$.
The key insight from the analysis in Section 3 is if the model has bias on a given dataset, this bias is likely to be more ubiquitous, hence evident, in the ill-defined versions of the tasks.
Let $A$ be an attack on the well-defined task that removes one of the task component $C_{i}$ that makes this task defined insufficiently, i.e. $A(T)=\hat{T}$ and $P(\hat{T})=0$ while initially $P(T)=1$. Let $A(D)$ be an attack on dataset $D$ which means applying element-wise the same attack $A$ to each $T$ from $D$. Subsequently, we consider three types of attacks that make the task ill-defined: $A_{v=0}$ for all zeroed frames in the video, $A_{q=0}$ for all questions as an empty string, and $A_{o=0}$ for all options as IDs only.
We investigate the importance of the following components of the bias $Bias_{M}$ : $Bias_{M}|_{A_{v=0}(D)}$, $Bias_{M}|_{A_{q=0}(D)}$, $Bias_{M}|_{A_{o=0}(D)}$, and how efficiently we can handle the general bias by balancing through these components.
@@ 4.2.2 Debiasing
First, we estimate the model bias on the dataset after attack A. We derive the $P_{p}$ knowing that $P_{d}$ is uniformly distributed from Equation  1:
@@
To estimate the bias under $A_{v=0}$, $A_{q=0}$ and $A_{o=0}$ attacks, we first compute $P_{p}(d_{i}|A_{v=0}(T))$, $P_{p}(d_{i}|A_{q=0}(T))$ and $P_{p}(d_{i}|A_{o=0})$, and then the prior estimation, respectively: $\tilde{P}_{p}(d_{i}|T)=softmax[P_{p}(d_{i}|A_{v=0}(T))+P_{p}(d_{i}|A_{q=0}(T))+P_{p}(d_{i}|A_{o=0}(T))]$.
@@
To gain the optimal weights, we use a 5-fold cross-validation procedure. In each fold, we optimize weights $\{w_{i}\}$ using the COBYLA algorithm Powell (1994), subject to the constraints $0\leq w_{i}\leq 1$ or $|w_{i}|\leq 1$.
The restriction $0\leq w_{i}\leq 1$ implies that the resulting debiasing vector is formed as a convex combination of the prior directions, remaining within the positive span of the original vectors. With this easy-to-interpret setup, we analyze how much each prior contributes to the final debiasing vector. Notably, a weight distribution with $w_{i}=0$ enables ablation-style analysis by turning off the corresponding debiasing component. Uninformative or detrimental priors are automatically suppressed, resulting in a more streamlined and potentially more generalizable debiasing vector. When we allow the more relaxed constraint $|w_{i}|\leq 1$, the COBYLA optimizer explores the full linear subspace spanned by the prior vectors, and the solution is no longer restricted to the positive orthant. This provides additional flexibility to correct for complex biases that may not be adequately addressed by additive combinations alone.
While we minimize the standard deviation of Recall_std across the options on the test set, we monitor the same bias metrics on the validation set. After completing the cross-validation we average the estimated global priors across the folds and apply this averaged global prior. Both algorithms, BOLD and Weighted_BOLD, are outlined in Appendix D.
@@ 5 Results
Both BOLD and Weighted_BOLD consistently improve performance metrics for Video-LLaMA and Video-LLaVA. SeViLA generally shows more resilience to bias calibration, with minimal effect on performance on NExT-QA and marginal deterioration on STAR. This may be attributed to the model’s training for statistically close to uniform distribution of results over a fixed options number. Despite decreases in Accuracy and F1 Mean on STAR, the reduction in Recall_std across all datasets for this model suggests that its debiased responses are more truthful, reflecting a fairer distribution of correct answers across options.
@@ 6 Conclusion
We presented the first comprehensive analysis of selection bias in video-based MCQA tasks, using modified datasets to pinpoint how bias emerges within specific answer options and across different VLM architectures. Our findings reveal that while all models share common tendencies in answer distribution, they also possess distinct bias profiles both across and within datasets. Notably, the bias vector is consistently more pronounced when projecting onto the decomposed planes where one key MCQA component (video, question, or answer options) is removed.
Building on these insights, we proposed the BOLD calibration method using a global prior bias vector, derived from the decomposed versions of the task, to mitigate selection bias. Our technique requires only three decompositions per question and is thus resource-efficient. Our BOLD bias removal not only enhances fairness and interpretability but also improves performance metrics. Extending the method to Weighted_BOLD refines the balance between debiasing and performance gains.
This easily scalable, post-processing, automatic, and unsupervised method is applicable to any existing VLM and allows for a sharper focus on its true reasoning capabilities. By making MCQA benchmarks more robust, transparent, and impartial, we enable a clearer assessment of the genuine reasoning cues in VLMs and pave the way for future improvements in model design and training strategies.
Furthermore, elevating bias suppression to an abstract level, where tasks are systematically decomposed and key components are removed to identify bias, we offer a framework beyond the current scope to address other bias-related challenges in machine learning models.
@@ 7 Limitations
Our approach rests on the assumption that when models lack sufficient direct information, they rely on indirect cues, manifesting as selection bias. While we have shown that bias can be decomposed by selectively removing the video, question, or answer component, this decomposition may not be exhaustive. There could be more optimal directions in the latent space of decomposed priors. Similarly, although our method identifies pronounced bias when all answers are correct, there remain scenarios where further supervised debiasing (e.g., applying COBYLA optimization with true labels) might yield more optimal results. We aimed for an easy-to-implement, unsupervised solution without relying on ground-truth answers.
We also assume the validity of uniform distribution as a baseline: Jensen-Shannon Divergence operates under the assumption that, in the absence of prior information, the distribution should be uniform. If the test MCQA data or task suggest an uneven distribution of answer options, this may lead to underestimating the significance of certain factors.
Another limitation is that we primarily assessed our post-processing calibration on the original datasets rather than systematically verifying performance improvements across all modifications. While our current evaluation covers four challenging video datasets and three model architectures, more granular, scenario-specific testing would provide stronger evidence of robustness. However, performing comprehensive evaluations on every modification poses practical challenges. These limitations point toward future work: exploring more nuanced testing scenarios, investigating methods that do not rely on uniform reference distributions, and potentially leveraging additional computational resources or innovative techniques for even more thorough bias and performance assessments.
@@ 8 Ethics Statement
We based our experiments on existing datasets, ensuring compliance with their respective copyrights. Our research uses open-source VLMs, which, while offering accessibility and flexibility, also carry the inherent risks associated with open-ended text generation.
@@ Acknowledgements
Olga Loginova thanks Amazon Alexa for supporting her research through a generous donation to Raffaella Bernardi. Ravi Shekhar has been partially funded by the UKRI and the European Union. The views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. All authors are deeply grateful to Giovanni Bonetta for his invaluable contributions and insightful consultations at earlier stages of the project.
@@ A.1 NExT-QA and NExT-GQA
NExT-QA, released in 2021, is one of the first datasets designed to assess a VLM’s ability to capture causal, temporal, and descriptive topics through what/when/why/how and before/after questions. It comprises videos sourced from the VidOR dataset Shang et al. (2019), with an average length of 35.73 seconds. The test split primarily contains causal questions (52.57%) addressing why and how, temporal reasoning questions (31.03%) about preceeding, following, or concurrent actions, and descriptive tasks (16.04%). Videos are manually annotated with questions and correct answers. Distractors are selected from 50 candidates based on a cosine similarity of less than 0.9 to the correct answer, ensuring they are close but not identical. This approach sometimes allows questions to be meaningfully answered without viewing the video (see Figures  68).
@@
NExT-GQA, released in 2024 as an extension of NExT-QA, focuses on evaluating the localization abilities of VLMs. It includes manually annotated timestamp labels that are key for answering the questions. Descriptive questions are excluded because they pertain to global video content or have answers available throughout the video with no specific temporal grounding. We also excluded QA pairs with multiple annotated time intervals. Consequently, the test set of NExT-GQA comprises 56.69% causal questions and 43.31% temporal questions.
@@ A.2 STAR
The question-answer pairs are generated using functional programs based on situation hyper-graphs: the graphs represent the extraction of abstract representations from the videos, while the functional programs provide multiple types of questions that cover different levels of difficulty in situated reasoning. There are four types of questions focusing on occurred facts, temporal order, future probability, and feasibility in specific situations:
@@ A.3 Perception Test
We use a 40% split of the correct-answer-annotated validation set to save the resources. Sampling is stratified to maintain the original balance of question types and the correct answer distribution. The split contains approximately 15% closed questions, which poses an additional challenge for models that tend to exhibit affirmation bias Pezeshkpour and Hruschka (2023). Notably, for closed questions in this dataset, Video-LLaMA and Video-LLaVA often respond with "yes" or "no" rather than selecting an answer option.
@@ A.4 Video-MME
Video-MME was developed to address the lack of diversity in video MCQA benchmarks. It consists of 900 YouTube videos ranging from 11 seconds to 1 hour in length. The dataset covers 6 primary visual domains with 30 subfields and 12 task types including questions on temporal and spatial perception, object and action reasoning, OCR, and counting problems. All QA pairs and distractors are manually created with three questions per video. Additionally, the dataset features videos in languages other than English. The videos additionally are supplied with subtitles.
The released test split has an imbalanced option distribution, so we equalized it to properly assess option bias.
@@ B.1 Video-LLaMA
Video-LLaMA aligns a BLIP-2-based visual encoder Li et al. (2023) with LLaMA Touvron et al. (2023). It incorporates audio signals using Imagebind Girdhar et al. (2023). Both the video and audio Q-formers are trained on the Webvid-2M dataset Bain et al. (2021) to align the modalities. LLaMA is fine-tuned with visual instructional data.
@@
To overcome this, we experimented with various prompts, temperature settings, and token limits, and developed the following strategy:
@@ B.2 Video-LLaVA
Video-LLaVA (Lin et al., 2023a) uses the pre-trained CLIP ViT-L/14 vision encoder Radford et al. (2021) and the Vicuna language model Chiang et al. (2023). Visual features from the encoder are projected into the same-dimensional space as the textual embeddings via a linear projector.
Video-LLaVA effectively delivers concise answers as just a single option based on a short, general prompt. However, in some cases, the model outputs as an answer option only "A" or responds with "yes" or "no" to a closed question.
@@ B.3 SeViLA
SeViLA adopts a selfchaining approach with BLIP-2 Li et al. (2023) for keyframe localization and answering question. Its Localizer uniformly samples 32 frames and selects the 4 most relevant key ones, while the Answerer generates responses based on those frames. The Localizer and Answerer both have Flan-T5 Chung et al. (2022) for LLM.
@@
Appendix C Models’ Performance on All Settings
C.1 Experimental Settings Examples
As the modifications to the video component do not affect the questions and answers, Table 4 presents examples of adjustments applied only to the question-answer pairs, along with their original versions. The example provided for the Video-MME dataset illustrates analogous changes implemented across all four datasets.
C.2 Models’ Performance on the Settings
Tables 5 through 16 present the distribution of selections across answer options and the accuracy (where applicable) for all settings, along with the target distributions. Notably, as soon as the correct answer aligns with the model’s biased option, the performance boosts.
The full data on distribution answers in each option is given in Figures 25 to 27.
Appendix D Debiasing Algorithms and Results
We fixed the random seed to 1 and conducted a series of experiments varying the weights and the sampling parameter $k$. All results including the default values are presented in Tables  1720.
Table 4: Questions and answers modifications example from MME dataset. Correct answer is highlighted in bold in the “Default” modification type.
Modification Type Question Answers
Default How many national flags appear in the video? (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Empty answers Default question (a0) – (a1) – (a2) – (a3) –
Answer shuffling Default question (a0) 4. (a1) 5. (a2) 2. (a3) 3.
Additional empty option Default question (a0) 3. (a1) 5. (a2) 2. (a3) 4. (a4) –
Correct answer in each option Default question (a0) 4. (a1) 4. (a2) 4. (a3) 4.
All identical answers (a0) Default question (a0) 3. (a1) 3. (a2) 3. (a3) 3.
All identical answers (a1) Default question (a0) 5. (a1) 5. (a2) 5. (a3) 5.
All identical answers (a2) Default question (a0) 2. (a1) 2. (a2) 2. (a3) 2.
All identical answers (a3) Default question (a0) 4. (a1) 4. (a2) 4. (a3) 4.
Correct answer in (a0) Default question (a0) 4. (a1) 5. (a2) 2. (a3) 3.
Correct answer in (a1) Default question (a0) 3. (a1) 4. (a2) 2. (a3) 5.
Correct answer in (a2) Default question (a0) 3. (a1) 5. (a2) 4. (a3) 2.
Correct answer in (a3) Default question (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Correct answer (a0) with shuffling Default question (a0) 4. (a1) 5. (a2) 3. (a3) 2.
Correct answer (a1) with shuffling Default question (a0) 5. (a1) 4. (a2) 2. (a3) 3.
Correct answer (a2) with shuffling Default question (a0) 3. (a1) 2. (a2) 4. (a3) 5.
Correct answer (a3) with shuffling Default question (a0) 5. (a1) 3. (a2) 2. (a3) 4.
Rephrased question What is the total number of national flags that appear in the video? (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Empty questions (a0) 3. (a1) 5. (a2) 2. (a3) 4.